Find the most general antiderivative for the function 6x−−√3−5x2−−√3.
Antiderivative =
1
Expert's answer
2020-12-02T17:35:46-0500

Mathematically, anti-derivative function is given by; ∫xn= (xn+1)/ (n+1) (+C)

Therefore, we proceed to calculate the antiderivative of 6x−−√3−5x2−−√3

First of all, we simplify our equation as follows;

(6x)--(√3)−(5x2)--(√3) = 6x + √3 - 5x2 + √3 = 6x + (√3 + √3) - 5x2 = 6x + 2√3 - 5x2 = -5x2 + 6x + 2√3.

we then procced to integrate function -5x2 + 6x + 2√3.

(-5x2+1/2+1) + (6x1+1/1+1) + (2√3x0+1/0+1) + C

(-5x3/3) + (6x2/2) + (2√3x1/1) + C= (-5/3)x3 + (6/2)x2 + (2√3x1/1) + C

therefore, the general antiderivative = (-5/3)x3 + 3x2 + 2√3x + C.